OpenStudy (anonymous):

To convert a Fahrenheit temperature to Celsius, one formula to use is F=9/5C+32, where F is the Fahrenheit temperature (in degrees) and C is the Celsius temperature. What is the Celsius temperature (to the nearest degree) when Fahrenheit temperature is 77 degrees?

OpenStudy (anonymous):

\[77=\frac{ 9 }{ 5 }C+32\] \[77-32=\frac{ 9 }{ 5 }C+32-32\] \[45=\frac{ 9 }{ 5}C\] \[\frac{ 45 }{ \frac{ 9 }{ 5 } }=\frac{ \frac{ 9 }{ 5 } }{ \frac{ 9 }{ 5 } }C\] 25=C

OpenStudy (anonymous):

Thank you very much!!!